# Diffusion Model Based Signal Recovery Under 1-Bit Quantization

### v1

@@ Introduction
Experimental setup. Our experiments employ pretrained DMs that correspond to image datasets at a resolution of $\times256$. Specifically, for FFHQ dataset (karras2019style), we utilized the score model (song2020improved) for QCS-SGM and the denoising model provided by (chung2023diffusion) for Diff-OneBit, DPS, DAPS, DiffPIR and SIM-DMIS. For CelebA (karras2018progressive) datasets, we used the models pretrained by (ho2020denoising). [^1] Since QCS-SGM does not provide a configuration file for reproducing results on the CelebA dataset, and we have not yet trained a score model on CelebA, we only tested the performance of QCS-SGM on the FFHQ dataset. For ImageNet (deng2009imagenet), we utilized the models pretrained by (dhariwal2021diffusion). All the experiments are conducted on NVIDIA Tesla V100.
@@
Baselines and metrics. We compare our proposed Diff-OneBit approach with several state-of-the-art methods. The primary baselines are specialized solvers for quantized measurements, including QCS-SGM and SIM-DMIS. Additionally, we include DiffPIR, DPS, and DAPS, that can handle non-linear measurement models, in our comparison. However, their non-linear frameworks are incompatible with 1-bit quantization, as the link function is non-differentiable or cannot be explicitly expressed. Thus, for our comparison, we use their linear versions, ignoring the non-linear link function. Reconstruction quality is evaluated using a comprehensive set of metrics. For fidelity and faithfulness to the ground truth, we use the peak signal-to-noise ratio (PSNR) and the structural similarity index measure (SSIM). For perceptual quality, we use the Learned Perceptual Image Patch Similarity (LPIPS) (zhang2018unreasonable) score and the Fréchet Inception Distance (FID) (heusel2017gans).
@@
Table 1: Quantitative results of 1-bit CS on FFHQ, CelebA and ImageNet images. We compare our method with DiffPIR, DPS, DAPS QCS-SGM and SIM-DMIS. Input images have an additive Gaussian noise of $\sigma=0.5$ and $\sigma=1.5$.
@@
Table 2: Quantitative results of 1-bit CS on FFHQ images under different NFEs. We evaluate our method under 20, 50 and 100 NFEs with additive Gaussian noise of $\sigma=0.5$.
Quantitative results
For 1-bit CS, the main quantitative results for the FFHQ, CelebA and ImageNet datasets are presented in Table 1. Our proposed method, Diff-OneBit, consistently outperforms all baseline methods across all evaluation metrics, demonstrating its superior reconstruction accuracy. For logistic regression, the main quantitative results on the FFHQ dataset are presented in Table 3. We compare our method with DPS, DAPS, and DiffPIR, which are capable of handling non-linear measurements. We also compare with SIM-DMIS, as logistic regression is a special case of SIM.
Our experiments demonstrate that existing solvers for linear or non-linear measurement models perform poorly in 1-bit quantization tasks. Tables 1 and 3 show that DiffPIR, DPS, and DAPS significantly underperforms compared to specialized methods and our proposed Diff-OneBit approach. These results demonstrate that optimization and gradient based methods, designed for linear or non-linear problems, are ineffective for 1-bit measurement models.
@@
Table 3: Quantitative results of logistic regression on FFHQ images. We compare Diff-OneBit with DAPS, DPS, DiffPIR, and SIM-DMIS without additional Gaussian noise, setting NFE to 20 for Diff-OneBit, 100 for DiffPIR, 150 for SIM-DMIS, and 1000 for DPS and DAPS.
Qualitative results
@@
Ablation studies and further analysis
@@
Experiment under varying NFEs. We evaluate our method by varying NFEs from 20 to 100 for 1-bit CS on 100 FFHQ images with Gaussian noise $\sigma=0.5$. Table 2 demonstrates that our method is efficient even at low NFEs. Moreover, the reconstruction quality further improves as NFE increases from 20 to 100, showing the scalability of our method.
@@ Conclusion
Appendix A Supplementary Material
Diffusion Model Based Signal Recovery under 1-Bit Quantization
Experimental results on 1-bit random inpainting
To further validate the effectiveness of our proposed method on 1-bit quantization tasks, we introduce 1-bit random inpainting. The degradation model for this masked image inpainting task can be expressed as:
@@ Appendix A Supplementary Material
We compare Diff-OneBit against DDRM, DDNM+, and DiffPIR, which can handle noisy linear measurement models. For DDRM, DDNM+, and DiffPIR, We use linear measurements which contain substantially more information than the 1-bit quantized measurements that our method uses. We test 100 images on FFHQ validation sets. The results summarized in Table 4 and visualized in Figure 8, clearly demonstrate the superiority of our approach.
@@
Table 4: Quantitative results of 1-bit random inpainting on FFHQ images. We consider random inpainting with a 50% mask ratio and an additive Gaussian noise $\sigma=0.5$. We compare Diff-OneBit with DDRM, DDNM+ and DiffPIR, setting NFE=20 for all methods.
Quantitative analysis. As shown in Table 4, Diff-OneBit achieves significantly better quantitative metrics. It obtains a PSNR of $20.89$, an SSIM of $0.61$, and an LPIPS of $0.38$. Remarkably, our method outperforms DiffPIR and DDNM+ even though it operates on severely degraded, 1-bit measurements. For example, the best-performing linear baseline, DDRM, achieves an LPIPS of $0.40$ with linear measurement, whereas our method achieves lower LPIPS with 1-bit measurement. This highlights the effectiveness of our proposed method on 1-bit inpainting random.
@@
The influence of the Gaussian noise in Diff-OneBit
@@
1:Measurements $\mathbf{y}\in\mathbb{R}^{M}$, degradation matrix $\mathbf{A}\in\mathbb{R}^{M\times N}$, pre-trained noise prediction network $\bm{\epsilon_{\theta}}$, maximum number of iterations $K$, data-fidelity function $\mathcal{L}(\cdot;\mathbf{y})$, schedules $\{\alpha_{t_{k}},\sigma_{t_{k}}\}_{k=0}^{K}$, penalty coefficient $\lambda$, pre-calculated $\mu_{k}=\lambda(\alpha_{t_{k}}^{2}/\sigma_{t_{k}}^{2})$, stochasticity level $\zeta$
@@
5:  $\mathbf{\tilde{z}}_{t_{k}}\leftarrow(\mathbf{\tilde{x}}_{t_{k}}-\sigma_{t_{k}}\bm{\epsilon}_{\theta}(\mathbf{\tilde{x}}_{t_{k}},t_{k}))/\alpha_{t_{k}}$
@@
To investigate this, we conducted experiments on the FFHQ validation sets using 100 images for the 1-bit CS task with an additive Gaussian noise level of $\sigma=0.5$. For these experiments, the penalty coefficient $\lambda$ was set to 0.05, and we varied $\zeta$ across the set $\{0,0.2,0.5,0.8,1.0\}$. The quantitative results are presented in Table 5.
@@
Table 5: The performance of Diff-Onebit under different stochasticity level $\zeta$ for 1-bit CS on FFHQ images with $\sigma=0.5$. We set $\lambda=0.05$ and NFE=20 for all the experiments.
Ablation study on hyperparameters $\lambda$ and $\sigma$
In this section, we present evaluation results for hyperparameters of our method. Regarding $\lambda$, we conducted an ablation study with values of 0.01, 0.02, 0.05, and 0.1 on 1-bit CS using FFHQ images. Regarding $\sigma$, we evaluated 1-bit CS using FFHQ images with values of 0.5, 1.0, 1.5, 2.0, and 2.5. The quantitative results are summarized in Table 6 and Table 7, respectively.
Table 6: Quantitative results of Diff-OneBit on FFHQ dataset under different $\lambda$ values for 1-bit CS.
@@
Table 7: Quantitative results of Diff-OneBit on FFHQ dataset under different $\sigma$ values for 1-bit CS.
@@
As shown in Table 6, $\lambda=0.05$ achieves the best performance across all metrics, demonstrating an optimal balance between data consistency and regularization. Table 7 indicates that smaller $\sigma$ values lead to better reconstruction quality, with $\sigma=1.0$ yielding the highest PSNR and SSIM scores while maintaining the lowest LPIPS.
Computational efficiency
@@
As presented in Table 8, SIM-DMIS is the quickest method as expected. SIM-DMIS utilizes measurements in the beginning of an unconditional sampling process, thus avoiding costly iterative guidance. Our method, Diff-OneBit, ranks as the second, significantly outperforming other iterative guidance methods like DPS, DAPS, and DiffPIR. The efficiency of Diff-OneBit stems from its low NFE. However, each NFE involves 100 inner iteration steps to solve the optimization subproblem, which contributes to the overall runtime. This trade-off allows for strong guidance with minimal NFEs. The results clearly highlight that our approach offers a compelling balance, providing a much more efficient solution than traditional guided methods.
@@
Table 8: Comparison of inference time (in seconds) for reconstructing 10 images on FFHQ. We compare Diff-OneBit with DPS, DAPS, DIffPIR, SIM-DMIS and QCS-SGM for 1-bit CS on FFHQ images.
Implementation details of our method and baselines
@@
DDRM. For 1-bit inpainting, we adopt the recommended values from (kawar2022denoising), specifically $\eta=0.85$ and $\eta_{b}=1.0$.
DDNM. Similarly, for 1-bit inpainting, we follow the recommended value from (wang2023zero), setting $\eta=0.85$.
DPS. We find that the learning rate $\zeta$ recommended by (chung2023diffusion), i.e., $\zeta=0.3$, is too large for compressed sensing tasks. Therefore, we use a constant value $\zeta=0.1$ for both 1-bit CS and logistic regression.
DAPS. Following the recommended configuration from DAPS-1K in (zhang2025improving), we set 5 ODE steps and 200 annealing steps. Additionally, we perform Langevin dynamics for 100 iterations per timestep with step size $\eta_{t}=\eta_{0}(\delta+t/T(1-\delta)),$ where $\eta_{0}=1\times 10^{-4}$, $\delta=0.01$, and $T=1000$.
DiffPIR. We employ the recommended values from (zhu2023denoising), namely the DDIM hyperparameter $\eta=1.0$ and $\lambda=7.0$. Furthermore, we set the number of optimization steps to 50 and the learning rate to 0.1.
QCS-SGM. For 1-bit CS, we adopt the recommended values from (meng2023quantized).
SIM-DMIS. We observe that the recommended values from (tang2025learning) are too small for $3\times 256\times 256$ images. Consequently, we adjust the values from $C_{s}=C_{s}^{\prime}=55$ to $C_{s}=C_{s}^{\prime}=135$.
@@
Out-of-distribution evaluation of diffusion priors
To evaluate the generalization capability of diffusion priors across different data distributions, we conduct out-of-distribution (OOD) experiments. Specifically, we test our Diff-OneBit method on FFHQ validation images using a DM trained on the CelebA dataset, rather than on the FFHQ dataset. This setup assesses whether the learned diffusion prior can effectively regularize inverse problems even when the test distribution differs from the training distribution. In practice, we perform 1-bit CS reconstruction with Gaussian noise $\sigma=0.5$ on 100 FFHQ validation images, utilizing a DM that was pre-trained on the CelebA dataset. Table 9 presents the quantitative results of OOD evaluation, demonstrating that the diffusion prior learned from CelebA transfers effectively to FFHQ images, despite the distribution shift between the two datasets.
@@
Table 9: Out-of-distribution evaluation of diffusion priors. We compare the reconstruction performance on FFHQ validation images using DMs trained on FFHQ versus CelebA. The marginal performance gap demonstrates the generalization capability of diffusion priors across different image distributions.
From MAP objective to alternating optimization with diffusion prior
@@
$\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]=\frac{\mathbf{x}_{t}-\sigma_{t}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)}{\alpha_{t}},$ (44)
where $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ is the noise prediction from the pre-trained DM. Since $\tilde{\mathbf{x}}_{t}=\hat{\mathbf{x}}^{(k+1)}$, the corresponding unscaled observation is $\mathbf{x}_{t}=\alpha_{t}\hat{\mathbf{x}}^{(k+1)}$. Substituting this into Tweedie’s formula yields the z-update:
$\hat{\mathbf{z}}^{(k+1)}=\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}=\alpha_{t}\hat{\mathbf{x}}^{(k+1)}]=\frac{\alpha_{t}\hat{\mathbf{x}}^{(k+1)}-\sigma_{t}\bm{\epsilon}_{\theta}(\alpha_{t}\hat{\mathbf{x}}^{(k+1)},t)}{\alpha_{t}}.$ (45)
Eq. (45) establishes the relationship between the clean image estimate $\hat{\mathbf{z}}^{(k+1)}$ and the unscaled observation $\mathbf{x}_{t}=\alpha_{t}\hat{\mathbf{x}}^{(k+1)}$. The predicted clean image $\hat{\mathbf{z}}^{(k+1)}$ inherently incorporates the learned image prior from the DM through the noise prediction network $\bm{\epsilon}_{\theta}$, and serves as the prior-regularized estimate for the next x-update. As the penalty parameter $\mu$ increases through the reverse diffusion process, the constraint $\mathbf{x}=\mathbf{z}$ is increasingly enforced, and the alternating scheme converges to a solution that balances data fidelity with the diffusion prior, thereby solving the original MAP estimation problem in Eq. (12).
Norm estimation
Under 1-bit quantization, it is typically difficult to estimate the norm of the underlying signal (knudson2016one). To demonstrate the capability of Diff-OneBit for norm estimation, we evaluated Diff-OneBit across varying levels of Gaussian noise in 1-bit CS and logistic regression on the FFHQ dataset. We tested pre-quantization noise levels at $\sigma\in\{0.5,1.0,1.5,2.0,2.5\}$ for 1-bit CS and $\sigma\in\{0,0.5,1.0,1.5,2.0\}$ for logistic regression. In logistic regression, Diff-OneBit is designed for the noiseless scenario with $\sigma=0$, and we simply reused this method for norm estimation. As shown in Figure 11, our method generally keeps the relative error between the $\ell_{2}$-norms of the recovered vector $\tilde{\mathbf{x}}$ and the underlying signal $\mathbf{x}^{*}$, defined as ${|\|\tilde{\mathbf{x}}\|_{2}-\|\mathbf{x}^{*}\|_{2}|}/{\|\mathbf{x}^{*}\|_{2}}$, within a small range.

### v2

@@ IV Experiments
Experimental setup. Our experiments employ pretrained DMs that correspond to image datasets at a resolution of $\times256$. Specifically, for FFHQ dataset [32], we utilized the score model [62] for QCS-SGM and the denoising model provided by [11] for Diff-OneBit, DPS, DAPS, DiffPIR and SIM-DMIS. For CelebA [30] datasets, we used the models pretrained by [24]. [^1] Since QCS-SGM does not provide a configuration file for reproducing results on the CelebA dataset, and we have not yet trained a score model on CelebA, we only tested the performance of QCS-SGM on the FFHQ dataset. For ImageNet [15], we utilized the models pretrained by [16]. All the experiments are conducted on NVIDIA Tesla V100.
@@
Baselines and metrics. We compare our proposed Diff-OneBit approach with several state-of-the-art methods. The primary baselines are specialized solvers for quantized measurements, including QCS-SGM and SIM-DMIS. Additionally, we include DiffPIR, DPS, and DAPS, that can handle non-linear measurement models, in our comparison. However, their non-linear frameworks are incompatible with 1-bit quantization, as the link function is non-differentiable or cannot be explicitly expressed. Thus, for our comparison, we use their linear versions, ignoring the non-linear link function. Reconstruction quality is evaluated using a comprehensive set of metrics. For fidelity and faithfulness to the ground truth, we use the peak signal-to-noise ratio (PSNR) and the structural similarity index measure (SSIM). For perceptual quality, we use the Learned Perceptual Image Patch Similarity (LPIPS) [73] score and the Fréchet Inception Distance (FID) [23].
@@
Table I: Quantitative results of 1-bit CS on FFHQ, CelebA and ImageNet images. We compare our method with DiffPIR, DPS, DAPS QCS-SGM and SIM-DMIS. Input images have an additive Gaussian noise of $\sigma=0.5$ and $\sigma=1.5$.
@@
Table II: Quantitative results of 1-bit CS on FFHQ images under different NFEs. We evaluate our method under 20, 50 and 100 NFEs with additive Gaussian noise of $\sigma=0.5$.
IV-A Quantitative results
For 1-bit CS, the main quantitative results for the FFHQ, CelebA and ImageNet datasets are presented in Table I. Our proposed method, Diff-OneBit, consistently outperforms all baseline methods across all evaluation metrics, demonstrating its superior reconstruction accuracy. For logistic regression, the main quantitative results on the FFHQ dataset are presented in Table III. We compare our method with DPS, DAPS, and DiffPIR, which are capable of handling non-linear measurements. We also compare with SIM-DMIS, as logistic regression is a special case of SIM.
Our experiments demonstrate that existing solvers for linear or non-linear measurement models perform poorly in 1-bit quantization tasks. Tables I and III show that DiffPIR, DPS, and DAPS significantly underperforms compared to specialized methods and our proposed Diff-OneBit approach. These results demonstrate that optimization and gradient based methods, designed for linear or non-linear problems, are ineffective for 1-bit measurement models.
@@ IV-A Quantitative results
Table III: Quantitative results of logistic regression on FFHQ images. We compare Diff-OneBit with DAPS, DPS, DiffPIR, and SIM-DMIS without additional Gaussian noise, setting NFE to 20 for Diff-OneBit, 100 for DiffPIR, 150 for SIM-DMIS, and 1000 for DPS and DAPS.
IV-B Qualitative results
@@
IV-C Ablation studies and further analysis
@@
Experiment under varying NFEs. We evaluate our method by varying NFEs from 20 to 100 for 1-bit CS on 100 FFHQ images with Gaussian noise $\sigma=0.5$. Table II demonstrates that our method is efficient even at low NFEs. Moreover, the reconstruction quality further improves as NFE increases from 20 to 100, showing the scalability of our method.
@@ V Conclusion
Appendix A Experimental results on 1-bit random inpainting
To further validate the effectiveness of our proposed method on 1-bit quantization tasks, we introduce 1-bit random inpainting. The degradation model for this masked image inpainting task can be expressed as
@@ Appendix A Experimental results on 1-bit random inpainting
We compare Diff-OneBit against DDRM, DDNM+, and DiffPIR, which can handle noisy linear measurement models. For DDRM, DDNM+, and DiffPIR, We use linear measurements which contain substantially more information than the 1-bit quantized measurements that our method uses. We test 100 images on FFHQ validation sets. The results summarized in Table IV and visualized in Figure 8, clearly demonstrate the superiority of our approach.
@@
Table IV: Quantitative results of 1-bit random inpainting on FFHQ images. We consider random inpainting with a 50% mask ratio and an additive Gaussian noise $\sigma=0.5$. We compare Diff-OneBit with DDRM, DDNM+ and DiffPIR, setting NFE=20 for all methods.
Quantitative analysis. As shown in Table IV, Diff-OneBit achieves significantly better quantitative metrics. It obtains a PSNR of $20.89$, an SSIM of $0.61$, and an LPIPS of $0.38$. Remarkably, our method outperforms DiffPIR and DDNM+ even though it operates on severely degraded, 1-bit measurements. For example, the best-performing linear baseline, DDRM, achieves an LPIPS of $0.40$ with linear measurement, whereas our method achieves lower LPIPS with 1-bit measurement. This highlights the effectiveness of our proposed method on 1-bit inpainting random.
@@
Appendix B The influence of the Gaussian noise in Diff-OneBit
@@
1:Measurements $\mathbf{y}\in\mathbb{R}^{M}$, degradation matrix $\mathbf{A}\in\mathbb{R}^{M\times N}$, pre-trained noise prediction network $\bm{\epsilon_{\bm{\theta}}}$, maximum number of iterations $K$, data-fidelity function $\mathcal{L}(\cdot;\mathbf{y})$, schedules $\{\alpha_{t_{k}},\sigma_{t_{k}}\}_{k=0}^{K}$, penalty coefficient $\lambda$, pre-calculated $\mu_{k}=\lambda(\alpha_{t_{k}}^{2}/\sigma_{t_{k}}^{2})$, stochasticity level $\zeta$
@@
5:  $\mathbf{\tilde{z}}_{t_{k}}\leftarrow(\mathbf{\tilde{x}}_{t_{k}}-\sigma_{t_{k}}\bm{\epsilon}_{\bm{\theta}}(\mathbf{\tilde{x}}_{t_{k}},t_{k}))/\alpha_{t_{k}}$
@@
To investigate this, we conducted experiments on the FFHQ validation sets using 100 images for the 1-bit CS task with an additive Gaussian noise level of $\sigma=0.5$. For these experiments, the penalty coefficient $\lambda$ was set to 0.05, and we varied $\zeta$ across the set $\{0,0.2,0.5,0.8,1.0\}$. The quantitative results are presented in Table V.
@@
Table V: The performance of Diff-Onebit under different stochasticity level $\zeta$ for 1-bit CS on FFHQ images with $\sigma=0.5$. We set $\lambda=0.05$ and NFE=20 for all the experiments.
Appendix C Ablation study on hyperparameters $\lambda$ and $\sigma$
In this section, we present evaluation results for hyperparameters of our method. Regarding $\lambda$, we conducted an ablation study with values of 0.01, 0.02, 0.05, and 0.1 on 1-bit CS using FFHQ images. Regarding $\sigma$, we evaluated 1-bit CS using FFHQ images with values of 0.5, 1.0, 1.5, 2.0, and 2.5. The quantitative results are summarized in Table VI and Table VII, respectively.
@@ Appendix C Ablation study on hyperparameters $\lambda$ and $\sigma$
Table VI: Quantitative results of Diff-OneBit on FFHQ dataset under different $\lambda$ values for 1-bit CS.
@@
Table VII: Quantitative results of Diff-OneBit on FFHQ dataset under different $\sigma$ values for 1-bit CS.
As shown in Table VI, $\lambda=0.05$ achieves the best performance across all metrics, demonstrating an optimal balance between data consistency and regularization. Table VII indicates that smaller $\sigma$ values lead to better reconstruction quality, with $\sigma=1.0$ yielding the highest PSNR and SSIM scores while maintaining the lowest LPIPS.
Appendix D Computational efficiency
@@
As presented in Table VIII, SIM-DMIS is the most efficient method as expected. SIM-DMIS utilizes measurements in the beginning of an unconditional sampling process, thus avoiding costly iterative guidance. Our method, Diff-OneBit, ranks as the second, significantly outperforming other iterative guidance methods like DPS, DAPS, and DiffPIR. The efficiency of Diff-OneBit stems from its low NFE. However, each NFE involves 100 inner iteration steps to solve the optimization subproblem, which contributes to the overall runtime. This trade-off allows for strong guidance with minimal NFEs. The results clearly highlight that our approach offers a compelling balance, providing a much more efficient solution than traditional guided methods.
@@
Table VIII: Comparison of inference time (in seconds) for reconstructing 10 images on FFHQ. We compare Diff-OneBit with DPS, DAPS, DIffPIR, SIM-DMIS and QCS-SGM for 1-bit CS on FFHQ images.
Appendix E Implementation details of our method and baselines
@@
DDRM. For 1-bit inpainting, we adopt the recommended values from [33], specifically $\eta=0.85$ and $\eta_{b}=1.0$.
DDNM. Similarly, for 1-bit inpainting, we follow the recommended value from [68], setting $\eta=0.85$.
DPS. We find that the learning rate $\zeta$ recommended by [11], i.e., $\zeta=0.3$, is too large for compressed sensing tasks. Therefore, we use a constant value $\zeta=0.1$ for both 1-bit CS and logistic regression.
DAPS. Following the recommended configuration from DAPS-1K in [72], we set 5 ODE steps and 200 annealing steps. Additionally, we perform Langevin dynamics for 100 iterations per timestep with step size $\eta_{t}=\eta_{0}(\delta+t/T(1-\delta)),$ where $\eta_{0}=1\times 10^{-4}$, $\delta=0.01$, and $T=1000$.
DiffPIR. We employ the recommended values from [75], namely the DDIM hyperparameter $\eta=1.0$ and $\lambda=7.0$. Furthermore, we set the number of optimization steps to 50 and the learning rate to 0.1.
QCS-SGM. For 1-bit CS, we adopt the recommended values from [47].
SIM-DMIS. We observe that the recommended values from [65] are too small for $3\times 256\times 256$ images. Consequently, we adjust the values from $C_{s}=C_{s}^{\prime}=55$ to $C_{s}=C_{s}^{\prime}=135$.
@@
Appendix F Out-of-distribution evaluation of diffusion priors
To evaluate the generalization capability of diffusion priors across different data distributions, we conduct out-of-distribution (OOD) experiments. Specifically, we test our Diff-OneBit method on FFHQ validation images using a DM trained on the CelebA dataset, rather than on the FFHQ dataset. This setup assesses whether the learned diffusion prior can effectively regularize inverse problems even when the test distribution differs from the training distribution. In practice, we perform 1-bit CS reconstruction with Gaussian noise $\sigma=0.5$ on 100 FFHQ validation images, utilizing a DM that was pre-trained on the CelebA dataset. Table IX presents the quantitative results of OOD evaluation, demonstrating that the diffusion prior learned from CelebA transfers effectively to FFHQ images, despite the distribution shift between the two datasets.
@@ Appendix F Out-of-distribution evaluation of diffusion priors
Table IX: Out-of-distribution evaluation of diffusion priors. We compare the reconstruction performance on FFHQ validation images using DMs trained on FFHQ versus CelebA. The marginal performance gap demonstrates the generalization capability of diffusion priors across different image distributions.
Appendix G From MAP objective to alternating optimization with diffusion prior
@@
$\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]=\frac{\mathbf{x}_{t}-\sigma_{t}\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)}{\alpha_{t}},$ (45)
where $\bm{\epsilon}_{\bm{\theta}}(\mathbf{x}_{t},t)$ is the noise prediction from the pre-trained DM. Since $\tilde{\mathbf{x}}_{t}=\hat{\mathbf{x}}^{(k+1)}$, the corresponding unscaled observation is $\mathbf{x}_{t}=\alpha_{t}\hat{\mathbf{x}}^{(k+1)}$. Substituting this into Tweedie’s formula yields the z-update:
$\hat{\mathbf{z}}^{(k+1)}=\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}=\alpha_{t}\hat{\mathbf{x}}^{(k+1)}]=\frac{\alpha_{t}\hat{\mathbf{x}}^{(k+1)}-\sigma_{t}\bm{\epsilon}_{\bm{\theta}}(\alpha_{t}\hat{\mathbf{x}}^{(k+1)},t)}{\alpha_{t}}.$ (46)
Eq. (46) establishes the relationship between the clean image estimate $\hat{\mathbf{z}}^{(k+1)}$ and the unscaled observation $\mathbf{x}_{t}=\alpha_{t}\hat{\mathbf{x}}^{(k+1)}$. The predicted clean image $\hat{\mathbf{z}}^{(k+1)}$ inherently incorporates the learned image prior from the DM through the noise prediction network $\bm{\epsilon}_{\bm{\theta}}$, and serves as the prior-regularized estimate for the next x-update. As the penalty parameter $\mu$ increases through the reverse diffusion process, the constraint $\mathbf{x}=\mathbf{z}$ is increasingly enforced, and the alternating scheme converges to a solution that balances data fidelity with the diffusion prior, thereby solving the original MAP estimation problem in Eq. (13).
Appendix H Norm estimation
Under 1-bit quantization, it is typically difficult to estimate the norm of the underlying signal [35]. To demonstrate the capability of Diff-OneBit for norm estimation, we evaluated Diff-OneBit across varying levels of Gaussian noise in 1-bit CS and logistic regression on the FFHQ dataset. We tested pre-quantization noise levels at $\sigma\in\{0.5,1.0,1.5,2.0,2.5\}$ for 1-bit CS and $\sigma\in\{0,0.5,1.0,1.5,2.0\}$ for logistic regression. In logistic regression, Diff-OneBit is designed for the noiseless scenario with $\sigma=0$, and we simply reused this method for norm estimation. As shown in Figure 11, our method generally keeps the relative error between the $\ell_{2}$-norms of the recovered vector $\tilde{\mathbf{x}}$ and the underlying signal $\mathbf{x}^{*}$, defined as ${|\|\tilde{\mathbf{x}}\|_{2}-\|\mathbf{x}^{*}\|_{2}|}/{\|\mathbf{x}^{*}\|_{2}}$, within a small range.